A conducting sphere of radius R and carrying a charge Q is joined to an uncharged conducting sphere of radius 2R, the charge transferred from one sphere to the other will be...

Grade:

1 Answers

Anurag Kishore
14 years ago

Hi Cathy

On joining the two spheres their potential will become equal

Let the charge remaining on charged sphere be q then by charge conservation, charge on uncharged sphere will be (Q - q).

Their potential will be equal so,

Kq/R = K(Q-q)/2R

=>  q = Q/3

So charge transferred from initially charged sphere to initially uncharged sphere is Q - Q/3 = 2Q/3